# EARL: Entropy-Aware RL Alignment of LLMs for Reliable RTL Code Generation

### v1

@@ Abstract
Recent advances in large language models (LLMs) have demonstrated significant potential in hardware design automation, particularly in using natural language to synthesize Register-Transfer Level (RTL) code. Despite this progress, a gap remains between model capability and the demands of real-world RTL design, including syntax errors, functional hallucinations, and weak alignment to designer intent. Reinforcement Learning with Verifiable Rewards (RLVR) offers a promising approach to bridge this gap, as hardware provides executable and formally checkable signals that can be used to further align model outputs with design intent. However, in long, structured RTL code sequences, not all tokens contribute equally to functional correctness, and naïvely spreading gradients across all tokens dilutes learning signals. A key insight from our entropy analysis in RTL generation is that only a small fraction of tokens (e.g., always, if, assign, posedge) exhibit high uncertainty and largely influence control flow and module structure. To address these challenges, we present EARL, an Entropy-Aware Reinforcement Learning framework for Verilog generation. EARL performs policy optimization using verifiable reward signals and introduces entropy-guided selective updates that gate policy gradients to high-entropy tokens. This approach preserves training stability and concentrates gradient updates on functionally important regions of code. Our experiments on VerilogEval and RTLLM show that EARL improves functional pass rates over prior LLM baselines by up to 14.7%, while reducing unnecessary updates and improving training stability. These results indicate that focusing RL on critical, high-uncertainty tokens enables more reliable and targeted policy improvement for structured RTL code generation.
I Introduction
@@
Large Language Models (LLMs) have achieved remarkable success in natural language processing and general-purpose code generation, demonstrating strong capabilities in understanding context, modeling semantics, and synthesizing coherent outputs across a wide range of tasks [2024llmse_survey, 2025llmcode_survey]. While these models have revolutionized software development workflows, applying LLMs to specialized domains such as hardware design remains a challenging frontier [2025llmeda_survey].
@@
One such challenge is the generation of Register-Transfer Level (RTL) code, typically written in Verilog. RTL code lies as the foundation for digital hardware design, translating architectural intent into synthesizable logic. Writing RTL code requires deep domain knowledge and significant engineering effort. Small errors in syntax, port declarations, or timing semantics can lead to non-compilable or functionally incorrect designs. Automating RTL code generation has the potential to significantly accelerate hardware development cycles, reduce engineering effort, and make hardware design more accessible. Recent efforts have explored applying LLMs to Verilog generation, with promising results in syntactic fluency and basic design translation [nyu2022benchmark, chipnemo, 2023chipgpt].
@@
However, achieving reliable RTL generation remains challenging. Most existing approaches rely on supervised fine-tuning (SFT) on pre-collected datasets [pku2024origen, xie2024rtlcoder]. While this improves syntactic fluency, compiler and testbench feedback, which is essential for functional correctness, remains unused during training. This motivates reinforcement learning with verifiable rewards (RLVR) [2025rl_verilog], which leverages executable and formally checkable signals to align model outputs with design intent. Yet, applying RLVR to RTL code generation introduces two fundamental difficulties. First, compiler and testbench feedback is inherently sparse and delayed, making credit assignment particularly challenging in long, structured code sequences [2025beyond]. Second, standard RLVR methods spread updates uniformly across all tokens, even though only a small subset governs structural and functional correctness. These limitations highlight the need for entropy-aware reinforcement learning methods that selectively emphasize critical tokens while preserving stability on deterministic ones.
Figure 1: Token-level entropy distribution in RTL generation.
@@
Addressing these challenges requires understanding how learning signals distribute across tokens in RTL sequences. To this end, we conduct an entropy analysis of RTL code generation, which further highlights the limitations of uniform updates. As shown in Fig. 1, the entropy distribution is highly skewed. Most tokens are generated with near-zero entropy, while only a small fraction exhibit significantly higher uncertainty. This pattern echoes recent findings in RLVR research [2025entropy_review, 2025beyond, 2025stabilizing]. In RTL, the imbalance is even more pronounced, since deterministic syntax and port declarations dominate sequences, while high-entropy tokens correspond to module structure, control flow, and signal connections that largely determine correctness. These findings motivate the need for reinforcement learning updates that focus on critical, high-entropy tokens rather than treating all positions uniformly. We provide a more detailed entropy analysis in RTL generation in Section III.
@@
To address these challenges, we propose EARL, an Entropy-Aware Reinforcement Learning framework for Verilog generation. EARL builds on supervised initialization and utlizes verifiable reward signals, e.g., syntax validity, interface compliance, and functional correctness. Crucially, EARL introduces entropy-guided selective updates. Instead of spreading gradients uniformly across all tokens, it concentrates policy updates on high-entropy tokens that disproportionately determine module structure, control flow, and downstream correctness. By concentrating policy gradients on uncertain and functionally critical tokens, EARL improves alignment with hardware correctness while preserving training stability. Our contributions are summarized as follows:
  1. 1.
@@
  2. 2.
    Entropy-aware RL framework. We introduce EARL, an entropy-aware RL framework that integrates verifiable rewards with selective policy updates on high-entropy tokens. Unlike uniform update schemes, EARL targets high-entropy tokens while preserving coherence on low-entropy ones, improving alignment with hardware correctness and stabilizing training dynamics. The design is optimizer-agnostic and can be layered on standard objectives. We instantiate it with both PPO and DAPO in this work.
  3. 3.
@@
II Background
II-A Large Language Model for Verilog Code Generation
Recent progress in large language models (LLMs) has spurred research into thier application to Verilog code generation [nyu2022benchmark, xie2024rtlcoder, nyu2024, 2024autovcoder, ren2024rtlfixer, 2024assertllm, pku2024origen, 2024betterv, haven2025]. Existing approaches can be grouped into three categories: supervised fine-tuning, prompt engineering, and reinforcement learning.
Supervised fine-tuning (SFT) methods adapt pre-trained LLMs using hardware-specific datasets. Examples include RTLCoder [xie2024rtlcoder] and VeriGen [nyu2024], which fine-tune on automatically generated instruction–code pairs. Other works enrich data through domain-specific transformations or augmentations. For instance, BetterV [2024betterv] leverages Verilog-to-C translation. Despite their gains, SFT methods are constrained by limited and noisy datasets and remain static, lacking adaptive feedback.
Prompt engineering approaches guide models via specialized prompting or interactive tool calls. RTLFixer [ren2024rtlfixer] employs ReAct prompting and retrieval to iteratively fix syntax errors. OriGen [pku2024origen] combines exemplar prompting with self-reflection and code-to-code transformations, while HaVen [haven2025] introduces a hallucination taxonomy and employs chain-of-thought prompting over symbolic modalities. Prompt-based methods can improve syntactic and functional correctness, but often suffer from high computational cost and poor scalability due to reliance on repeated synthesis or tool feedback.
Reinforcement learning (RL) has recently been explored to bridge the gap between fluency and functional correctness. Building on RLHF [2020rlhf, 2022rlhf], methods like PPOCoder [2023ppocoder] and PLUM [2024plum] leverage compiler or test outcomes as reward signals. For Verilog generation, RLVR-style methods have emerged [2025rl_1, rl2025verireason], but existing designs are limited: rewards based on AST similarity can be gamed without functional benefit [2025rl_1], while structured reasoning with distillation may constrain diversity [rl2025verireason]. These limitations highlight the need for reinforcement learning frameworks that leverage verifiable multi-signal feedback while avoiding uniform or inefficient token-level updates.
@@ II-A Large Language Model for Verilog Code Generation
Figure 2: Word cloud visualization of token-level entropy in RTL generation: (a) high-entropy tokens and (b) low-entropy tokens.
II-B RLVR Algorithms
Proximal Policy Optimization (PPO). PPO [schulman2017proximal] is a widely used policy gradient algorithm that stabilizes reinforcement learning by clipping the probability ratio between new policy $\pi_{\theta}$ and old policy $\pi_{\theta_{\mathrm{old}}}$. The objective is as follows:
$J_{\mathrm{PPO}}(\theta)=$~ $\mathbb{E}_{(q,a)\sim\mathcal{D},\,o\sim\pi_{\theta_{\mathrm{old}}}(\cdot\mid q)}\Big[\min\!\big(r_{t}(\theta)\,\hat{A}_{t},\,$
@@ II-B RLVR Algorithms
where $\mathcal{D}$ is a dataset of queries $q$ and corresponding ground-truth answers $a$. $\mathbf{o}=(o_{1},\ldots,o_{T})$ denotes a sampled output sequence from the old policy $\pi_{\theta_{\mathrm{old}}}$, $\epsilon\in\mathbb{R}$ is a clipping hyperparameter, and $\hat{A}_{t}$ is the estimated advatage computed using a value network.
Group Relative Policy Optimization (GRPO). GRPO [grpo2024] is a variant of PPO designed to reduce the cost of critic training in large-scale LLM alignment. The objective is as follows:
$J_{\mathrm{GRPO}}(\theta)=$ $~\mathbb{E}_{(q,a)\sim\mathcal{D},\,\{o^{i}\}_{i=1}^{G}\sim\pi_{\theta_{\mathrm{old}}}(\cdot\mid q)}\Bigg[\frac{1}{\sum_{i=1}^{G}\lvert o^{i}\rvert}\sum_{i=1}^{G}\sum_{t=1}^{\lvert o^{i}\rvert}$
@@
where, instead of learning a separate value function, GRPO computes token-level advantages $\hat{A}_{t}^{i}$ by normalizing rewards $R^{i}$ within a group of $G$ sampled responses $\mathbf{o}=(o^{1},\ldots,o^{G})$ and $r^{i}_{t}(\theta)$ for each $o^{i}$:
$\hat{A}_{t}^{i}$ $=\frac{\,r^{i}-\operatorname{mean}\!\big(\{R^{i}\}_{i=1}^{G}\big)\,}{\operatorname{std}\!\big(\{R^{i}\}_{i=1}^{G}\big)},$
$\text{where}\quad R^{i}$ $=\begin{cases}1.0,&\text{if }\mathrm{is\_equivalent}\!\left(a,\,o^{i}_{t}\right),\\ 0.0,&\text{otherwise.}\end{cases}$
Dynamic Sampling Policy Optimization (DAPO). DAPO [2025dapo] extends GRPO via clip-higher, dynamic sampling, token-Level Policy Gradient Loss, and overlong reward shaping techiniques.
Entropy-aware RLVR. Recent studies have highlighted the importance of token-level entropy in reinforcement learning with verifiable rewards. Recent works in math reasoning tasks [2025beyond, 2025stabilizing, 2025entropy_review] have shown that high-entropy minority tokens act as critical decision points that determine downstream reasoning and structural correctness. In contrast, low-entropy tokens typically correspond to deterministic elements such as syntax completion or port declarations, contributing little to downstream reasoning outcome. These findings suggest that RL methods should focus gradient updates on high-entropy tokens, rather than uniformly updating all tokens. Inspired by these findings, we examine entropy patterns in RTL generation and design EARL to exploit high-entropy tokens in Verilog code.
Figure 3: Heatmap of a code snippet.
Figure 4: EARL: an Entropy-Aware Reinforcement Learning framework.
III Token-Level Entropy in RTL Code Generation
While prior studies have examined entropy in chain-of-thought reasoning [2025entropy_review, 2025beyond, 2025stabilizing], we analyze entropy in RTL code generation at token level. Specifically, the entropy of the $t$-th token $H_{t}$ is defined as:
$H_{t}$ $=-\sum_{v=1}^{V}p_{t,v}\,\log p_{t,v},$ (1)
@@ III Token-Level Entropy in RTL Code Generation
Here $V$ is the vocabulary size, $\mathbf{z}^{\,t}\in\mathbb{R}^{V}$ is the pre-softmax logits, and $T$ is the decoding temperature. Note that $H_{t}$ measures the uncertainty of the predictive distribution at the $t$-th token.
We use a DeepSeek-7B coder model fine-tuned on our curated dataset, generating over $2.1\times 10^{5}$ response tokens at decoding temperature $T=1.0$. For each token $t$, entropy $H_{t}$ is computed according to Eq. (1). Figure 1 presents the overall entropy distribution. The histogram is highly skewed: over 80% of tokens fall below $0.15$, while only a small fraction exceeds $0.6$, forming a heavy tail of high-uncertainty tokens. To identify what these extremes correspond to, Figure 2 shows the top-100 lowest-entropy tokens and top-100 highest-entropy tokens with frequency $\geq 10$. Furthermore, Figure 3 overlays entropy values onto an RTL code example, highlighting uncertainty concentrations at specific structural positions.
Two consistent patterns emerge:
(i) Deterministic bulk vs. uncertain critical tokens. Low-entropy tokens are dominated by predictable syntax or declarative elements (e.g., end, input, output, whitespace), reflecting Verilog grammar. High-entropy tokens cluster around decision-making constructs with examples as follows:
  • Process/sensitivity: always, posedge, negedge
  • Control flow: if, case, ternary ?:
  • Binding/connection: assign, bus indices [i]
  • Module heads: module, often followed by identifiers
(ii) Entropy concentrates at structural junctions. As shown in Fig. 3, the heatmap shows that high-entropy regions appear at sensitivity lists (posedge, negedge) and conditional branches, precisely where module behavior and timing diverge. This parallels the notion of “forking tokens” [2025beyond], but here the semantic forks represent hardware control flow and signal behavior, not natural language discourse.
@@
IV Entropy-Aware RL Framework
IV-A Overview of EARL

### v2

@@ Abstract.
Recent advances in large language models (LLMs) have demonstrated significant potential in hardware design automation, particularly in using natural language to synthesize Register-Transfer Level (RTL) code. Despite this progress, a gap remains between model capability and the demands of real-world RTL design, including syntax errors, functional hallucinations, and weak alignment to designer intent. Reinforcement Learning with Verifiable Rewards (RLVR) offers a promising approach to bridge this gap, as hardware provides executable and formally checkable signals that can be used to further align model outputs with design intent. However, in long, structured RTL code sequences, not all tokens contribute equally to functional correctness, and naïvely spreading gradients across all tokens dilutes learning signals. A key insight from our entropy analysis in RTL generation is that only a small fraction of tokens (e.g., always, if, assign, posedge) exhibit high uncertainty and largely influence control flow and module structure. To address these challenges, we present EARL, an Entropy-Aware Reinforcement Learning framework for Verilog generation. EARL performs policy optimization using verifiable reward signals and introduces entropy-guided selective updates that gate policy gradients to high-entropy tokens. This approach preserves training stability and concentrates gradient updates on functionally important regions of code. Our experiments on VerilogEval and RTLLM show that EARL improves functional pass rates over prior LLM baselines by up to 14.7%, while reducing unnecessary updates and improving training stability. These results indicate that focusing RL on critical, high-uncertainty tokens enables more reliable and targeted policy improvement for structured RTL code generation. We will release the code upon acceptance. An anonymized repository for review is available at https://anonymous.4open.science/r/EARL-1C25.
conference: The Chips to Systems Conference; July 26–29, 2026; Long Beach, CA
1. Introduction
@@
Large Language Models (LLMs) have achieved remarkable success in natural language processing and general-purpose code generation, demonstrating strong capabilities in understanding context, modeling semantics, and synthesizing coherent outputs across a wide range of tasks (Liu et al., 2024a; Jiang et al., 2025). While these models have revolutionized software development workflows, applying LLMs to specialized domains such as hardware design remains a challenging frontier (Pan et al., 2025).
@@
One such challenge is the generation of Register-Transfer Level (RTL) code, typically written in Verilog. RTL code lies as the foundation for digital hardware design, translating architectural intent into synthesizable logic. Writing RTL code requires deep domain knowledge and significant engineering effort. Small errors in syntax, port declarations, or timing semantics can lead to non-compilable or functionally incorrect designs. Automating RTL code generation has the potential to significantly accelerate hardware development cycles, reduce engineering effort, and make hardware design more accessible. Recent efforts have explored applying LLMs to Verilog generation, with promising results in syntactic fluency and basic design translation (Thakur et al., 2022; Liu et al., 2024b; Chang et al., 2023).
@@
However, achieving reliable RTL generation remains challenging. Most existing approaches rely on supervised fine-tuning (SFT) on pre-collected datasets (Cui et al., 2025; Liu et al., 2024c). While this improves syntactic fluency, compiler and testbench feedback, which is essential for functional correctness, remains unused during training. This motivates reinforcement learning with verifiable rewards (RLVR) (Wang et al., 2025b), which leverages executable and formally checkable signals to align model outputs with design intent. Yet, applying RLVR to RTL code generation introduces two fundamental difficulties. First, compiler and testbench feedback is inherently sparse and delayed, making credit assignment particularly challenging in long, structured code sequences (Wang et al., 2025d). Second, standard RLVR methods spread updates uniformly across all tokens, even though only a small subset governs structural and functional correctness.
Figure 1. Token-level entropy distribution in RTL generation.
@@
To understand where learning signals should be allocated, we analyze token-level entropy in RTL generation. As shown in Figure 1, the entropy distribution is highly skewed. Most tokens are generated with near-zero entropy, while only a small fraction exhibit significantly higher uncertainty. This pattern echoes recent findings in RLVR research (Deng et al., 2025; Wang et al., 2025d, a). In RTL, the imbalance is even more pronounced, since deterministic syntax and port declarations dominate sequences. These observations suggest that uniform updates waste gradient budget, motivating reinforcement learning (RL) methods that selectively focus on high-entropy, high-impact positions. A more detailed entropy study is presented in Section 3.
@@
To address these challenges, we propose EARL, an Entropy-Aware Reinforcement Learning framework for Verilog generation. EARL combines supervised initialization with verifiable multi-signal rewards (syntax, interface, functionality) and introduces entropy-guided selective updates. By concentrating policy gradients on uncertain and functionally critical tokens, EARL improves alignment with hardware correctness while preserving training stability. Our contributions are summarized as follows:
  1. (1)
@@
  2. (2)
    Entropy-aware reinforcement Learning framework. We introduce EARL, an entropy-aware RL framework that integrates verifiable rewards with selective policy updates on high-entropy tokens. Unlike uniform update schemes, EARL targets high-entropy tokens while preserving coherence on low-entropy ones, improving alignment with hardware correctness and stabilizing training dynamics. The design is optimizer-agnostic and can be layered on standard objectives. We instantiate it with both Proximal Policy Optimization (PPO) and Dynamic Sampling Policy Optimization (DAPO) in this work.
  3. (3)
@@
2. Background
2.1. LLM for Verilog Code Generation
Recent progress in LLMs has spurred research into thier application to Verilog code generation (Thakur et al., 2022; Liu et al., 2024c; Thakur et al., 2024; Gao et al., 2024; Tsai et al., 2024; Yan et al., 2025; Cui et al., 2025; Pei et al., 2024; Yang et al., 2025). Existing approaches can be grouped into three categories: supervised fine-tuning, prompt engineering, and reinforcement learning.
Supervised fine-tuning (SFT) methods adapt pre-trained LLMs using hardware-specific datasets. Examples include RTLCoder (Liu et al., 2024c) and VeriGen (Thakur et al., 2024), which fine-tune on automatically generated instruction–code pairs. Other works enrich data through domain-specific transformations or augmentations. For instance, BetterV (Pei et al., 2024) leverages Verilog-to-C translation. Despite their gains, SFT methods are constrained by limited and noisy datasets and remain static, lacking adaptive feedback.
Prompt engineering approaches guide models via specialized prompting or interactive tool calls. RTLFixer (Tsai et al., 2024) employs ReAct prompting and retrieval to iteratively fix syntax errors. OriGen (Cui et al., 2025) combines exemplar prompting with self-reflection and code-to-code transformations, while HaVen (Yang et al., 2025) introduces a hallucination taxonomy and employs chain-of-thought prompting over symbolic modalities. Prompt-based methods can improve syntactic and functional correctness, but often suffer from high computational cost and poor scalability due to reliance on repeated synthesis or tool feedback.
Reinforcement learning has recently been explored to bridge the gap between fluency and functional correctness. Building on RLHF (Stiennon et al., 2020; Ouyang et al., 2022), methods like PPOCoder (Shojaee et al., 2023) and PLUM (Zhang et al., 2024) leverage compiler or test outcomes as reward signals. For Verilog generation, RLVR-style methods have emerged (Wang et al., 2025c, e), but existing designs are limited: rewards based on AST similarity can be gamed without functional benefit (Wang et al., 2025c), while structured reasoning with distillation may constrain diversity (Wang et al., 2025e). These limitations highlight the need for reinforcement learning frameworks that leverage verifiable multi-signal feedback while avoiding uniform or inefficient token-level updates.
@@ 2.1. LLM for Verilog Code Generation
Figure 2. Word cloud visualization of token-level entropy in RTL generation. Font size indicates the frequency of a lexical token across the corpus, while color encodes its average entropy (blue = low, orange = high).
2.2. RLVR Algorithms
Proximal Policy Optimization (PPO). PPO (Schulman et al., 2017) is a widely used policy gradient algorithm that stabilizes reinforcement learning by clipping the probability ratio between new policy $\pi_{\theta}$ and old policy $\pi_{\theta_{\mathrm{old}}}$. The objective is as follows:
(1) $J_{\mathrm{PPO}}(\theta)=$ $\mathbb{E}_{(q,a)\sim\mathcal{D},\,\tilde{o}\sim\pi_{\theta_{\mathrm{old}}}(\cdot\mid q)}\Big[\min\!\big(r_{t}(\theta)\,\hat{A}_{t},\,$
@@ 2.2. RLVR Algorithms
where $\mathcal{D}$ is a dataset of queries $q$ and corresponding ground-truth answers $a$. $\tilde{o}=(o_{1},\ldots,o_{T})$ denotes a sampled output sequence from the old policy $\pi_{\theta_{\mathrm{old}}}$, $\epsilon\in\mathbb{R}$ is a clipping hyperparameter, and $\hat{A}_{t}$ is the advantage estimate from a learned value function.
Group Relative Policy Optimization (GRPO). GRPO (Shao et al., 2024) removes the value function and computes the advantages by normalizing scalar rewards within a group of rollouts sampled from the same prompt. This group-wise normalization removes the need for a critic, reducing memory cost and improving stability. The objective is as follows:
(2) $J_{\mathrm{GRPO}}(\theta)=$ $~\mathbb{E}_{(q,a)\sim\mathcal{D},\,\{\tilde{o}^{i}\}_{i=1}^{G}\sim\pi_{\theta_{\mathrm{old}}}(\cdot\mid q)}\Bigg[\frac{1}{\sum_{i=1}^{G}\lvert\tilde{o}^{i}\rvert}\sum_{i=1}^{G}\sum_{t=1}^{\lvert\tilde{o}^{i}\rvert}$
@@
where each group contains $G$ rollouts $\{\tilde{o}^{i}\}_{i=1}^{G}$ sampled from the same query $q$. The advantage $\hat{A}_{t}^{i}$ is derived by normalizing rewards $R^{i}$ within a group of $G$ sampled responses $(\tilde{o}^{1},\ldots,\tilde{o}^{G})$ for each $\tilde{o}^{i}$:
(3) $\hat{A}_{t}^{i}$ $=\frac{\,R^{i}-\operatorname{mean}\!\big(\{R^{i}\}_{i=1}^{G}\big)\,}{\operatorname{std}\!\big(\{R^{i}\}_{i=1}^{G}\big)}.$
Dynamic Sampling Policy Optimization (DAPO). DAPO (Yu et al., 2025) extends GRPO via clip-higher, dynamic sampling, token-Level Policy Gradient Loss, and overlong reward shaping techiniques.
Entropy-aware RLVR. Recent studies have highlighted the importance of token-level entropy in reinforcement learning with verifiable rewards. Recent works in math reasoning tasks (Wang et al., 2025d, a; Deng et al., 2025) have shown that high-entropy minority tokens act as critical decision points that determine the trajectory of downstream reasoning. In contrast, low-entropy tokens typically correspond to routine, highly predictable components of the solution, contributing little to downstream reasoning outcome. These findings suggest that RL methods should focus gradient updates on high-entropy tokens, rather than uniformly updating all tokens. Inspired by these findings, we examine entropy patterns in RTL generation and design EARL to exploit high-entropy tokens in Verilog code.
(a) JC_counter.
(b) PE_single.
Figure 3. Token-level entropy visualizations for two Verilog modules. Each token is shaded according to its positional entropy, computed over the model’s distribution at that generation step. Only high-entropy tokens (above the 80th percentile within each response) are visually highlighted.
Figure 4. Overview of the EARL framework for Verilog generation. The top panel presents the end-to-end pipeline. The bottom panel details the three core components: dataset preparation (yellow), entropy-aware reinforcement learning (green), and reward-guided exploration (red).
3. Token-Level Entropy in RTL Code Generation
While prior studies have examined entropy in chain-of-thought reasoning (Deng et al., 2025; Wang et al., 2025d, a), we analyze entropy in RTL code generation at token level. Specifically, the entropy of the $t$-th token $H_{t}$ is defined as:
(4) $H_{t}$ $=-\sum_{v=1}^{V}p_{t,v}\,\log p_{t,v},$
@@ 3. Token-Level Entropy in RTL Code Generation
Here $V$ is the vocabulary size, $\mathbf{z}^{\,t}\in\mathbb{R}^{V}$ is the pre-softmax logits, and $T$ is the decoding temperature. Note that entropy is computed per position, not per token type. For example, the same lexical token (e.g., begin) may appear with low entropy in a deterministic context or high entropy in an ambiguous structural context. Thus, entropy reflects uncertainty at the generation step, not an inherent uncertainty of a token type. This metric is used throughout our analysis and later in the EARL training objective.
To analyze entropy behavior in RTL tasks, we begin with concrete examples. We compute entropy using a DeepSeek-7B coder model that we fine-tuned on our curated specification–implementation dataset. All rollouts use temperature at temperature $T=1.0$. For each generated token $t$, we record its entropy $H_{t}$ following the definition in Eq. (4). Figure 3 visualizes token-level entropy for two RTL modules from the RTLLM v1.1 benchmark (Lu et al., 2024). For each generated response, only tokens whose entropy exceeds the 80th percentile of that response are highlighted; low-entropy tokens are shown without background shading. For instance, in the JC_counter example (Figure 3(a)), control-structure keywords (e.g., if, begin) show higher entropy, purely syntactic or declarative tokens (e.g., end, endmodule, numeric literals) exhibit near-zero entropy. Similarly, in the PE_single module (Figure 3(b)), the accumulation logic (e.g., acc) exhibits elevated uncertainty due to its arithmetic and dependency structure.
Beyond these qualitative examples, we analyze entropy statistically over more than $2.1\times 10^{5}$ generated tokens produced by the same fine-tuned model. Figure 1 presents the overall entropy distribution. The distribution is heavily skewed: over 80% of tokens fall below $0.15$, while only a small minority exceed $0.6$, forming a distinct heavy tail of high-uncertainty positions. To identify what these extremes correspond to, Figure 2 shows the top-100 lowest-entropy tokens and top-100 highest-entropy tokens with frequency $\geq 10$. Two consistent patterns emerge:
(i) Deterministic bulk vs. uncertain critical tokens. Low-entropy positions correspond almost exclusively to rigid, grammar-driven Verilog constructs such as module headers, port lists, and white space. High-entropy positions, in contrast, occur at tokens that govern control or timing behavior—e.g., sensitivity triggers, conditional guards, and data‑path operations—reflecting genuine semantic uncertainty rather than lexical variability.
(ii) Entropy concentrates at structural junctions. As shown in Figure 3, the heatmap shows that high-entropy regions appear at sensitivity lists and conditional branches, precisely where module behavior and timing diverge. This parallels the notion of “forking tokens” (Wang et al., 2025d), but here the semantic forks represent hardware control flow and signal behavior, not natural language discourse.
@@
4. Entropy-Aware RL Framework
4.1. Overview of EARL